OpenStudy (anonymous):

x y 2 12 3 8 4 6 10 125 4.)Is the data in the table a direct variation or an inverse variation? Write an equation to model the data in the table. can some one explain the steps to me so i may get this answer plz

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Do you know what direct and inverse variation is?

OpenStudy (anonymous):

no

OpenStudy (igreen):

Direct variation means that when you multiply something to 'x', you get 'y'.

OpenStudy (igreen):

Inverse variation means that when you divide something by 'x', you get 'y'.

OpenStudy (igreen):

So is something being divided here or multiplied?

OpenStudy (igreen):

What's 12 / 6? @nemoo112233

OpenStudy (anonymous):

2

OpenStudy (igreen):

Yep, and what's 8 / 3?

OpenStudy (anonymous):

2.6?

OpenStudy (igreen):

Yes, since 2 and 2.6 aren't the same, we know that this isn't a direct variation, so it must be an inverse variation. Check: 2 * 12 = 24 3 * 8 = 24 4 * 6 = 24 So it's an inverse variation. \(\Large\color{lime}\checkmark\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @nemoo112233